# I Introduction

### v1

@@ Abstract
High-resolution imaging has revolutionized materials science by offering detailed insights into the atomic structures of materials. Electron microscopy and spectroscopy rely on analysing backscattered and transmitted electrons as well as stimulated radiation emission to form structural and chemical maps. These signals contain information about the elastic and inelastic electron-scattering processes within the sample, including collective and single electron excitations such as plasmons, inter- and intraband transitions. In this study, ab initio and Monte Carlo simulations were performed to investigate the behaviour of high-energy primary and secondary electrons in scanning transmission experiments on CsPbBr3 nanosamples. CsPbBr3 is a perovskite material known for its high photoluminescence quantum yield, making it promising for applications in light-emitting devices and solar cells. This study explores and estimates the reflection and transmission of primary and secondary electrons based on their kinetic energy as well as sample thickness and work function. The spatial distribution and energy spectra of the secondary electrons are also examined and calculated to understand their generation depth and energy dynamics. These findings establish a theoretical framework for studying electron-material interactions and can aid in optimizing scanning microscopy techniques for imaging and characterizing advanced materials.
@@ I Introduction
In transmission electron microscopy (TEM), a beam of high-energy electrons (typically 80-300 keV) is passed through a very thin sample ( 100nm) and interference between elastically scattered electrons is exploited to form an image. This approach enables a high resolution imaging down to the atomic resolution ( 0.1 nm) and allows the precise visualisation of atomic arrangements. In a scanning electron microscope (SEM) a focused beam of electrons is raster scanned across a sample and nm resolution images of the surface are formed from the analysis of emitted secondary electrons (SE) and backscattered electrons (BSE). The resolution in a SEM is largely limited by the interaction volume of the detected radiation which, in most cases, precludes the need for incident electron probes smaller than 1 nm. Scanning transmission electron microscopy (STEM) combines elements of SEM and TEM. A finely focused electron beam is raster scanned across a thin sample ( 100nm), and elastically scattered electrons are detected to form an image. At high incident electron energy and for thin samples, the interaction volume for elastic scattering is small and the microscope resolution is primarily limited by the size of the incident electron probe, which with modern aberration-corrected optics can reach sub-Ångström dimensions. Together with other emissions such as Auger electrons which are used to characterise the surface elemental compositions and X-rays applied to determine bulk crystal structures, the SEM, TEM and STEM represent the main techniques to analyse chemical composition and structure [3].
@@
In high-resolution STEM and TEM experiments a projection image is acquired and it is challenging to obtain information on sample variation along the beam direction. By placing SE detectors both before and after the sample within a STEM instrument, the surface sensitivity of SE detection can be exploited to determine structure on the entrance and exit surfaces of a thin sample [4, 5], to some extent overcoming the projection limitation of the STEM.
@@
A recent review [6], has highlighted the applications for studying the dynamics and transient phenomena of surface reconstructions, exsolution of catalysts, lunar and planetary materials and mechanical properties of 2D thin films utilising the simultaneous SE-STEM imaging.
Furthermore, inelastic scattering events within the sample, including SE generation [7] are a primary cause of electron beam-induced sample damage and need to be well understood to help develop novel approaches to imaging beam-sensitive materials.
@@
Motivated by the potential for transmitted SE imaging described in the above works, we construct a theoretical framework to calculate SE generation and emission in transmission through an industrially relevant metal-halide perovskite sample.
@@
Any electron microscopy technique relies on electron scattering in the sample, which includes both elastic and inelastic processes. Both spectroscopic and SE imaging are products of inelastic scattering processes within the sample. These mechanisms are driven by energy loss phenomena, including single-particle electron excitations, collective atomic modes such as plasmons, electron-phonon and electron-polaron interactions, and Auger decay. Part of the energy loss contributes to the generation of a secondary electron cascade and, upon emission from the surface, to the secondary electron yield (SEY). Secondary electrons appear in the electron energy spectrum as a tail with energies below 50 eV. The emission of secondary electrons in SEM plays a decisive role in the formation of the image contrast [8] and can also lead to unwanted charging of the sample, for example, [9].
@@
Electron energy loss phenomena are influenced by two factors, namely the properties of the beam (i.e. kinetic energy, radiation type and charge state) and the material properties, in particular the electronic excitation spectrum. The latter can be described by the so-called energy loss function (ELF), which can be obtained either from optical experiments or from EELS using a single scattering spectrum in which the elastic peak and multiple scattering have been removed [10] or via a Reverse Monte Carlo (RMC) method [11]. The ELF is a unique property of the material and is independent of the beam. However, it is difficult to interpret this from the experiment alone since the uniqueness of the ELF is not guaranteed by the fulfilment of the sum rules, i.e. the EELS signal can lead to different ELFs that all fulfil the sum rules [12] and thus to different beam stopping properties. This emphasises the need for more precise theoretical descriptions and predictions of electron energy loss. In this context, ab initio calculations can also be used to determine the dielectric properties, especially in the low energy loss region.
@@
In this article, the scattering behaviour of high-energy primary (up to 50 keV) and secondary (up to 50 eV) electrons in STEM experiments on CsPbBr3 thin films is investigated theoretically using mixed first-principles calculations and Monte Carlo (MC) approach. In particular, the linear response time-dependent density functional theory (LR-TDDFT) is used to calculate the ELF, while the charge transport in solids is modelled by an MC routine. We have also investigated the reflection and transmission of primary and secondary electrons as a function of the kinetic energy of incidence and the sample thickness. Since electron microscopy (EM) samples are susceptible to beam-inducing charging, we also determined the effects of the work function (WF) on electron emission. Finally, we determined the SE spectra and spatial distribution as a function of the initial kinetic energy of the primary beam by calculating the depth of secondary electron generation.
@@
Caesium lead bromide, CsPbBr3, is an inorganic perovskite material that has attracted much attention in optoelectronics due to its high photoluminescence quantum yield (PLQY) with a sharp, narrow-band blue-green light emission [13, 14, 15, 16, 17, 18]. These properties make it a good candidate for industrial applications, e.g. in light-emitting devices (LEDs). As CsPbBr3 is a purely inorganic material, it is more stable under heat and less susceptible to degradation under harsh environmental conditions [19, 20, 21]. Although CsPbBr3 has a relatively large bandgap (about 2.3 eV [22]), it can be used in tandem solar cells alongside silicon or other materials to cover different parts of the solar spectrum and increase overall efficiency. In addition, CsPbBr3 exhibits the “hot carrier effect” [23, 24], in which excited electrons maintain high energy levels longer than in other materials, which has the potential to further increase the efficiency of solar cells.
@@ II.1 Energy Loss Function
The energy loss processes of electrons are described using Ritchie’s dielectric theory [25, 26]. This approach is based on the knowledge of the macroscopic dielectric function , which can be calculated from first-principles using linear-response time-dependent density functional theory (LR-TDDFT) or obtained from optical experiments.
The former allows the inclusion of momentum dispersion, which is necessary for the calculation of the stopping properties of materials, with the same computational effort as in zero-momentum transfer, while the latter approach must be extended to finite momentum by analytical dispersion laws, which are known to be valid only within certain limits [26]. We have considered the orthorhombic geometric structure of CsPbBr3 [27, 22] shown in Fig. 1.
The electronic structure in the ground state was relaxed using the generalised gradient approximation (GGA) of the exchange-correlation functional as implemented in QE, Elk and Questaal. We have checked the convergence of the DFT parameters given in the literature [27]. We point out that the DFT and LR-TDDFT calculations performed with QE, Elk and Questaal were carried out within the random-phase approximation neglecting local field effects (RPA-NLFE). We have converged our -mesh grid up to 212121 points. The ALDA and RPA-LFE LR-TDDFT calculations were performed with the turboEELS code [28]. In the LR-TDDFT, the susceptibility function can be obtained by solving the following Dyson-like equation:
where q is the momentum transfer vector, is the absorbed energy, is the non-interacting (or independent) particle susceptibility calculated from the Kohn-Sham wave functions and is the Coulomb potential. In this work, we used the Adiabatic Local Density Approximation (ALDA) [29] to calculate the TDDFT kernel . The microscopic dielectric function of the material is related to the susceptibility as follows:
@@
If this relationship is applied to the momentum in the first Brillouin zone is referred to as the “optical limit”. The inversion of with G and G’ not equal to zero leads to the inclusion of the so-called local field effects (LFE), which take into account the inhomogeneity of the system, in contrast to the non-local field effect (NLFE), in which only the head of the matrix is inverted. The ELF can be written as:
To determine the effects of the ALDA correlation, the random phase approximation (RPA) to is also used within LFE and NLFE.
In the upper graph of Fig. (2) the ELFs are compared in the context of ALDA, RPA-LFE and RPA-NLFE. Since we are currently not able to compare our computer simulations with the experimental results, the peaks are convolved with a typical Gaussian broadening of 0.1 eV. When RPA-LFE is compared with RPA-NLFE, a dramatic change in the shape of the ELF can be observed due to the inhomogeneities. At the same time, the electron correlation in the ALDA kernel slightly increases the intensity in the 15-25 eV range, especially the 16.75 and 22 eV peaks, compared to RPA-LFE. The first less bright peak at 5 eV in the ALDA and RPA-LFE spectra is associated with the transition from the 4 states of the Br atom to the 6 states of Pb. The rich peak structures beyond 6 eV are primarily caused by transitions from the 4 states of the Br atom to a combination of 6 states of Cs and 6 states of Pb.
The imaginary and real parts of the dielectric matrix are shown in the bottom panel of Fig. (2). By analysing their spectral behaviour, it can be concluded that the 22 eV peak is associated with a mixed-state plasmon (the real part of the dielectric function remains positive, even if it is almost zero).
In Fig. (3), the TD-DFT ALDA ELF is merged with the experimental form factors from the NIST datasets [30]. This merging approach is due to the fact that high-energy TD-DFT ELF estimation is computationally too demanding due to the occurrence of numerous excited states. On the other hand, the extension from 100 eV to 400 keV, which is related to the atomic core-level excitations, is needed to accurately determine the total inelastic cross-section.
II.2 Total inelastic scattering cross-section and mean free path
@@
The equation for the inverse inelastic mean free path (IIMFP), from which the total inelastic scattering cross-section can be derived, is given by
where the lower and upper limits of the integral are obtained by the energy-momentum conservation laws, is the band gap, corresponds to the maximum energy of the incident electrons, is the Bohr radius and is the electron mass. Figure (4) shows the IMFP () for CsPbBr3 as a function of the electron kinetic energy. The IMFP minimum is reached at 5.34 Å for an incident electron energy of 57 eV.
II.3 Total elastic scattering cross-section and mean free path
Elastic scattering is accounted for in our MC approach using Mott theory, which we have generalised to account for the presence of multiple scattering between neighbours in the periodic unit cell [31]. In this approach, the differential elastic scattering cross-section (DESCS) for an electron scattered by a central atomic potential at an angle is as follows:
where and are the Mott amplitudes of the direct and spin-flip scattering for the atoms, which were determined using the relativistic partial wave expansion method [26] and indicates the distance between the th and th atom. and are the scattering and solid angles, respectively, measured in the laboratory frame. The total elastic scattering cross-section integrated over results as:
@@ III Monte Carlo results
Since we have access to the cross sections of elastic and inelastic scattering, our MC algorithm works as follows [32, 26]: We extract a random number that is sampled from a uniform distribution in the interval and use it to determine the step length that each electron travels within the solid target , where is the mean free path that includes all elastic () and inelastic () scattering processes and can be calculated from . Finally, the MC method selects an elastic or inelastic scattering event by comparing another random number , which is sampled from a uniform distribution in the range [0, 1], with the corresponding probability. If , it is an elastic scattering event that leads to an angular deviation of the electron’s trajectory, otherwise, the electron suffers an inelastic interaction that leads to both energy loss and (small) angular deviation [32]. We note that in order to escape from the surface of the sample, the electrons must overcome the barrier at the interface between the vacuum and the solid, i.e. they must fulfil the condition where is the angle formed between the surface normal and the intersection of the electron trajectory on the surface, and is the kinetic energy of the electrons. In addition, the transmission from solid to vacuum is modelled by a transmission coefficient that depends on the WF of the material [26].
III.1 Electron energy loss spectrum and secondary electron yield

### v2

@@ Abstract
High-resolution imaging has revolutionised materials science by providing detailed insights into the atomic structures of materials. Electron microscopy and spectroscopy rely on analysing backscattered and transmitted electrons as well as stimulated radiation emission to form structural and chemical maps. These signals contain information about the elastic and inelastic electron scattering processes within the sample, including collective and single electron excitations such as plasmons, inter- and intra-band transitions. In this study, ab initio and Monte Carlo simulations were performed to investigate the behaviour of high-energy primary and secondary electrons in scanning transmission experiments on CsPbBr3 nanosamples. CsPbBr3 is a perovskite material known for its high photoluminescence quantum yield, making it promising for applications in light-emitting devices and solar cells. This study investigates and estimates the reflection and transmission of primary and secondary electrons based on their kinetic energy, sample thickness and electron affinity. The spatial distribution and energy spectra of the secondary electrons are also analysed and calculated to understand their generation depth and energy dynamics. These findings provide a theoretical framework for the study of charge transport in perovskites and can help to optimise scanning microscopy techniques for the imaging and characterisation of advanced materials.
@@ I Introduction
In transmission electron microscopy (TEM), a beam of high-energy electrons is passed through a very thin sample ( 100 nm) and interference between elastically scattered electrons is exploited to form an image. This approach enables high-resolution imaging down to the atomic resolution ( 0.1 nm) and allows the precise visualisation of atomic arrangements. TEM microscopes typically operate at electron energies of 80-300 keV, although atomic resolution has already been demonstrated at 30 keV [3, 4] and even 15 keV [5].
In a scanning electron microscope (SEM), a focused beam of electrons is raster scanned across a sample, and nm resolution images of the surface are formed from the analysis of emitted secondary electrons (SE) and backscattered electrons (BSE). The resolution in a SEM is largely limited by the interaction volume of the detected radiation, which, in most cases, precludes the need for incident electron probes smaller than 1 nm. SEM microscopes typically operate at electron beam kinetic energies of 1-30 keV. Images in a SEM are generally formed by collecting secondary or backscattered electrons that originate from the top surface of the sample. Performing in a low voltage mode, SEM in transmission mode has advantages over TEM imaging (albeit with lower spatial resolution) due to the relatively low cost and simplicity of the SEM instrument [6]and has applications in both the physical [7] and biological [8] sciences.
@@
Scanning transmission electron microscopy (STEM) combines elements of SEM and TEM. A finely focused electron beam is raster scanned across a thin sample ( 100 nm), and scattered electrons are detected to form an image. At high incident electron energy and for thin samples, the interaction volume is small and the microscope resolution is primarily limited by the size of the incident electron probe, which, with modern aberration-corrected optics, can reach sub-Ångström dimensions. Together with other emissions such as Auger electrons which are used to characterise the surface elemental compositions and X-rays applied to determine bulk crystal structures, the SEM, TEM and STEM represent the main techniques to analyse chemical composition and structure of materials [9].
@@
In high-resolution STEM and TEM experiments, a projection image is acquired, and it is challenging to obtain information on a sample variation along the beam direction. By placing SE detectors both before and after the sample within a STEM instrument, the surface sensitivity of SE detection can be exploited to determine the structure on both entrance and exit surfaces of a thin sample [10, 11]. This approach, to some extent, overcomes the projection limitation of the STEM.
@@
A recent review [12] has highlighted the applications for studying the dynamics and transient phenomena of surface reconstructions, exsolution of catalysts, lunar and planetary materials and the mechanical properties of 2D thin films utilising the simultaneous SE-STEM imaging.
@@
Furthermore, inelastic scattering events within the sample, including SE generation, [13] are a primary cause of electron beam-induced sample damage and need to be well understood to develop new approaches for imaging beam-sensitive materials. Motivated by the potential for transmitted SE imaging described in the above works, we construct a theoretical framework to calculate SE generation and emission in transmission through an industrially relevant metal-halide perovskite sample.
@@
Any electron microscopy technique relies on electron scattering in the sample, which includes both elastic and inelastic processes. Both spectroscopic and SE imaging are products of inelastic scattering processes within the sample. These mechanisms are driven by energy loss phenomena, including single-particle electron excitations, collective atomic modes such as plasmons, electron-phonon and electron-polaron interactions, and Auger decay. Part of the energy loss contributes to the generation of a secondary electron cascade and, upon emission from the surface, to the secondary electron yield (SEY). Secondary electrons appear in the electron energy spectrum as a structured broad peak with energies below 50 eV. The emission of secondary electrons in SEM plays a decisive role in the formation of image contrast [14] and can also lead to unwanted charge of the sample [15].
@@
Electron energy loss phenomena are influenced by two factors, namely the properties of the beam (i.e. kinetic energy, radiation type, and charge state) and the material properties, in particular the electronic excitation spectrum. The latter can be described by the so-called energy loss function (ELF), which can be obtained either from optical experiments or from EELS using a single scattering spectrum with the elastic peak and multiple scattering removed [16] or via a Reverse Monte Carlo (RMC) method [17]. The ELF is a unique property of the material and is independent of the beam. However, it is difficult to interpret its spectral profile from the experiment alone, as the uniqueness of the ELF is not guaranteed by the fulfilment of the sum rules, i.e. the EELS signal can lead to different ELFs that all fulfil the sum rules [18] and thus to different beam stopping properties. This emphasises the need for more precise theoretical descriptions and predictions of electron energy loss. In this context, ab initio calculations can also be used to determine the dielectric properties [19], especially in the low energy loss region.
@@
In this article, the scattering behaviour of high-energy primary electrons (up to 50 keV, which covers the energy range of SEM and low energy TEM operation) and secondary electrons (up to 50 eV) in STEM experiments on CsPbBr3 thin films is theoretically investigated, using a mixed first-principles and Monte Carlo (MC) approach. In particular, linear response time-dependent density functional theory (LR-TDDFT) is used to calculate the ELF, while the MC routine models the charge transport in the solid. We have also investigated the reflection and transmission of primary and secondary electrons as a function of the kinetic energy of incidence and the sample thickness. Because electron microscopy (EM) samples are susceptible to beam-induced charging, we also determined the effects of the Electron Affinity (EA) on electron emission (we recall that the EA is measured from the bottom of the conduction band inside semiconductors and insulators). Finally, we determined the SE spectra and spatial distribution as functions of the initial kinetic energy of the primary beam by calculating the depth of secondary electron generation.
@@
Caesium lead bromide, CsPbBr3, is an inorganic perovskite material that has attracted much attention in optoelectronics due to its high photoluminescence quantum yield (PLQY) with a sharp, narrow-band blue-green light emission [20, 21, 22, 23, 24, 25]. These properties make it a good candidate for industrial applications, e.g. in light-emitting devices (LEDs). As CsPbBr3 is a purely inorganic material, it is more stable under heat and less susceptible to degradation under harsh environmental conditions [26, 27, 28]. Although CsPbBr3 has a relatively large bandgap (about 2.3 eV [29]), it can be used in tandem solar cells alongside silicon or other materials to cover different parts of the solar spectrum and increase overall efficiency. In addition, CsPbBr3 exhibits the “hot carrier effect” [30, 31], in which excited electrons maintain high energy levels longer than in other materials, which has the potential to further increase the efficiency of solar cells. It has also shown the possibility of trion generation under high-energy laser excitation [32].
@@ II.1 Energy Loss Function
The energy loss processes of electrons are described using Ritchie’s dielectric theory [33, 34]. This approach is based on the knowledge of the macroscopic dielectric function , which can be calculated from first principles using LR-TDDFT or obtained from optical experiments.
The former allows for the inclusion of momentum dispersion, which is necessary for the calculation of the stopping properties of materials, with the same computational effort as in zero-momentum transfer, whereas the latter approach must be extended to finite momentum by analytical dispersion laws, which are known to be valid only within certain limits [34]. We have considered the orthorhombic geometric structure of CsPbBr3 [35, 29], which belongs to the crystallographic space group and is shown in Fig. 1.
The electronic structure in the ground state was relaxed in the framework of density functional theory (DFT) using the generalised gradient approximation (GGA) of the exchange-correlation functional as implemented in the Quantum Espresso (QE) code suite [36]. We have checked the convergence of the DFT parameters given in the literature [37, 38, 35]. We note that the DFT and LR-TDDFT QE calculations performed under the random phase approximation (RPA) neglecting local field effects (RPA-NLFE) were compared with the same calculations using Elk [39], an all-electron full-potential linearised augmented plane-wave (LAPW) code and Questaal [40]. The RPA-LFE LR-TDDFT calculations were performed with the Liouville-Lanczos turboEELS code [41]. We have converged our -mesh grid to 212121 points.
In the LR-TDDFT, the susceptibility function can be obtained by solving the following Dyson-like equation:
where q is the momentum transfer vector, is the absorbed energy, is the non-interacting (or independent) particle susceptibility calculated from the Kohn-Sham wave functions and is the bare Coulomb potential. In this work, we used the Adiabatic Generalised Gradient Approximation (AGGA) [42] to calculate the TDDFT kernel . The microscopic dielectric function of the material is related to the susceptibility as follows:
@@
If this relationship is applied to the momentum in the first Brillouin zone is referred to as the “optical limit”. The inversion of with G and G’ not equal to zero leads to the inclusion of the so-called local field effects (LFE), which take into account the inhomogeneity of the system, in contrast to the non-local field effect (NLFE), in which only the head of the matrix is inverted. Using Eq. (3), the ELF can be written as:
To determine the effects of the AGGA correlation, the RPA () is also used within LFE and NLFE.
In the upper panel of Fig. 2 the ELFs are compared in the context of AGGA, RPA-LFE and RPA-NLFE. As we are currently unable to compare our computer simulations with the experimental results, the peaks are convolved with a typical Gaussian broadening of 0.1 eV.
When RPA-LFE is compared with RPA-NLFE, a dramatic change in the shape of the ELF can be observed due to inhomogeneities. At the same time, the electron correlation in the AGGA kernel slightly increases the intensity in the 15-25 eV range, especially the 16.75 and 22 eV peaks, compared to RPA-LFE. The first less bright peak at 5 eV in the AGGA and RPA-LFE spectra is associated with the transition from the 4 states of the Br atom to the 6 states of Pb. The rich peak structures beyond 6 eV are primarily caused by transitions from the 4 states of the Br atom to a combination of 6 states of Cs and 6 states of Pb.
The real and imaginary parts of the dielectric function calculated using LR-TDDFT within the AGGA framework are shown in the bottom panel of Fig. 2. By analysing their spectral behaviour, it can be concluded that the 22 eV peak is associated with a mixed-state plasmon (the real part of the dielectric function remains positive, even if it is almost zero).
In view of these results, we decided to use the TDDFT AGGA ELF, which contains the LFE, for our analysis.
In Fig. 3 the TDDFT AGGA LFE ELF is merged with the experimental form factors from the NIST datasets [43]. This merging is necessary because the estimation of the high-energy TD-DFT ELF is computationally too demanding because of the numerous excited states that must be included in the calculation. On the other hand, the extension from 100 eV to 400 keV, which is related to the atomic core-level excitations, is needed to accurately determine the total inelastic cross section.
II.2 Total inelastic scattering cross section and mean free path
@@
The equation for the inverse inelastic mean free path (IIMFP), from which the total inelastic scattering cross section can be derived, is given by
where the lower and upper limits of the integral are obtained by the energy-momentum conservation laws, is the Bohr radius, is the electron mass, and according to Ganachaud and Mokrani is the band gap and corresponds to the maximum energy of the incident electrons [44]. We have tested that the relativistic effects on the IMFP of CsPbBr3 are minimal by performing the calculations with the relativistic extension of Eq. 5 [45] (the maximum energy considered in this work is 1/10 of the rest mass of the electron). Fig. 4 shows the IMFP () for CsPbBr3 as a function of the electron kinetic energy.
II.3 Total elastic scattering cross section and mean free path
Elastic scattering is accounted for in our MC approach using Mott theory, which we have generalised to account for the presence of multiple scattering between neighbours in the periodic unit cell [46]. In this approach, the differential elastic scattering cross section (DESCS) for an electron scattered by a central atomic potential at an angle is as follows:
where and are the Mott amplitudes of direct and spin-flip scattering for the atoms, which were determined using the relativistic partial-wave expansion method [34] and the modulus of indicates the distance between the th and th atom. and are the scattering and solid angles, respectively, measured in the laboratory frame. The total elastic scattering cross section integrated over results as:
where .
@@ III Monte Carlo results
Since we have access to the cross sections of elastic and inelastic scattering, our MC algorithm works as follows [47, 34]: we extract a random number that is sampled from a uniform distribution in the interval and use it to determine the step length that each electron travels within the solid target , where is the mean free path that includes all elastic () and inelastic () scattering processes and can be calculated from .
Finally, the MC method selects an elastic or inelastic scattering event by comparing another random number , which is sampled from a uniform distribution in the range [0, 1], with the corresponding probability. If , it is an elastic scattering event that leads to an angular deviation of the electron’s trajectory; otherwise, the electron suffers an inelastic interaction that leads to both energy loss and (typically small) angular deviation [47]. We note that to escape from the surface of the sample, the electrons must overcome the barrier at the interface between the vacuum and the solid, i.e. they must fulfill the condition
where is the angle formed between the surface normal and the intersection of the electron trajectory on the surface, and is the kinetic energy of the electrons. In addition, the transmission from solid to vacuum is modeled by a transmission coefficient that depends on the EA of the material [34].
III.1 Electron energy loss spectrum
Using the MC approach described above, we have calculated the reflection electron energy loss spectrum (REELS) [48] for different initial energies of the incident electrons (black solid line for 10 keV kinetic energy, red line for 20 keV, green line for 30 keV, orange line for 40 keV, and blue line for 50 keV), as shown in Fig. 5.